OpenStudy (anonymous):

Solve the equation for the exact value of x. 2lnx=ln(6x+24)- ln3

OpenStudy (anonymous):

2lnx=ln(6x+24)- ln3 using the logarithm laws lnx^2=ln[(6x+24)/3] taking exponents on both sides e^lnx^2=e^ln[(6x+24)/3] x^2=[(6x+24)/3] x^2=2x+8 x^2-2x-8=0 factoring (x-4)(x+2)=0 x=4,-2 ans

OpenStudy (anonymous):

any question josh

OpenStudy (anonymous):

I get it. thanks!